OpenStudy (annipuppi):

use the binomial expression (p+q)^n to calculate a binomial distribution with n=5 and p=0.3. Please just post the steps in a simple, understandable way

OpenStudy (anonymous):

For a binomial distribution with parameters n = 5 , p = 0.3 . Find the probabilities of getting : (i) Atleast 3 successes. (ii) Atmost 3 successes. n = 5 , p = 0.3

OpenStudy (annipuppi):

I don't understand this at all... can you just list the steps with an explanation of each one. I usually get it better when I see all the steps together

OpenStudy (amistre64):

it looks like they asked their question in yours ....

OpenStudy (amistre64):

what do you understand about expanding a binomial?

OpenStudy (amistre64):

there is a long version, and a short version ... well, shorter at least. the formula for the expansion is much simpler to process than all of the distributing.

OpenStudy (amistre64):

p+q = 1 is something you need to know, since the sum of all probabilities is 1 (1)^n = 1 for all n, therefore (p+q)^n = 1 only if p+q = 1

OpenStudy (amistre64):

the formula for the expansion requires you to know factorials, combinatorics

OpenStudy (amistre64):

\[(p+q)^n=\sum_{k=0}^{n}\binom{n}{k}p^kq^{n-k}\] now the probability of x=k is defined by the kth term of the expansion \[P(x=k)=\binom{n}{k}p^{k}q^{n-k}\]